OpenStudy (anonymous):

help me with logs, log6 x = 1/4log6 16

OpenStudy (sasogeek):

ok so according to the law of logs \(\large log_ab^c=c(log_ab) \) this means that \(\large \frac{1}{4}log_616=log_616^{\frac{1}{4}}=log_62 \) so we dealt with the right side of the equation, let's put it together \(\huge log_6x=log_62 \) we can safely say that \(\huge x=2 \)